OpenStudy (anonymous):

Find the value of the constant b that makes the following function continuous on (-\infty,\infty). f(x) = 4x-5 if x < or = to 5 f(x) = -8x+b if x > 5

OpenStudy (anonymous):

The two functions will be continuous if they have the same value at x = 5, where the two functions switch off. So at x = 5, the first function: f(x) = 4x-5 f(5) = 4(5) - 5 = 15 So what you need to do is plug in 5 into the other function and find the value of b that will make the answer equal to 15.

OpenStudy (anonymous):

Okay, so 55 would be the answer. Thanks

OpenStudy (anonymous):

Yep :)